A Chemical reaction releases 350 kJ of energy. A joule (symbol J) is the unit of energy in the SI system of measurement. How much mass, in grams, was lost when this chemical reaction occurred? Hint: Use Einstein‘s equation E=mc^2. Use consistent SI units. If you substitute energy in joules and the velocity of light as 3.00*10^8 m/s, Your answer will come out in kilograms.
1
Expert's answer
2020-07-29T06:32:01-0400

350 kJ = 3.5*105 J

From Einstein's equation,

m = E / c2 = 3.5*105 J / (3.00*108 m/s)2 = 3.9*10-12 kg = 3.9*10-9 g

Answer: 3.9*10-9 g